OpenStudy (bellezabrilla):

Help with irrational square roots. (8th grade level)

OpenStudy (mathmale):

Please post the actual question you'd like help with. Thx.

OpenStudy (bellezabrilla):

\[4 \frac{ 5 }{ 11 } and \sqrt{20}\]

OpenStudy (bellezabrilla):

@mathmale okay so I'm given these pair of numbers and I have to write whether they are < , > , =

OpenStudy (freckles):

are you allowed to use a calculator? you can compare the decimal approximations of both numbers

OpenStudy (bellezabrilla):

No im not allowed to use a calculator @freckles

OpenStudy (freckles):

ok I have another idea

OpenStudy (freckles):

square both numbers

OpenStudy (freckles):

and compare the squares of both numbers

OpenStudy (anonymous):

\[4\frac{ 5 }{ 11 }=\frac{ 49 }{ 11 }=4.45454545....\]

OpenStudy (bellezabrilla):

OMG thank you so much !!!!!!!! @surjithayer

OpenStudy (anonymous):

yw